FEL Upgrade Project Weekly Brief
July 10-14, 2000


This week we ran the FEL for user tests. We had a very uneventful week with tests of nitriding a number on materials successfully performed. The processed surfaces show striking changes even to the naked eye. Work continues in trying to uncover the source of the beam asymmetry in the user labs.

Design/Construction Activities

Injector (WBS 4):
* The qualification bake-out of the Apertured Cesiator test system failed after the replacement vacuum pump also tripped at high current and became sensitive to mechanical impacts just like the first pump. A rigorous leak test showed no leaks. This atypical behavior for items constructed using ultra high vacuum techniques is a puzzle. We plan to find the source of the gas load by instrumentation and isolation.
* The first spare ceramic was successfully ion implanted at Berkeley and work on the second was intiated. This effort is proceeding well despite the two year hiatus in activity.

SRF (WBS 5):
* Tuner cost redesign has been completed and a prototype tuner is being procured.

* The BAFO for the helium vessels have been received from qualified vendors and are being reviewed.

* Space frame and vacuum vessel comments have been received from vendors and preparations for BAFO request will be completed and sent next week.

* Design work continues on remaining cryostat components; all are near completion.

RF Systems (WBS 6):

The HPEE group is making preparations for installing waveguide in the
3rd cryomodule position through the FEL ceiling penetrations during one
of the weekend shut downs during the upcoming user run.

Instrumentation (WBS 8)

* Completed prototype for optical BPM, testing with beam will begin next week. A quadrant LiTa detector is configured with preamps to give a scaled output of 1 volt = 1 mWatt of incident power. These signals will be connected with sample and hold circuits to provide an input to EPICS. EPICS will intern display a "Red October" plot of FEL position.

* Placed orders for beam viewer related items to support upgrade.

Beam Transport (WBS 9):

* Contrary to our original plan, two engineers from the engineering group were shifted to the SNS project. In future months, this will mothball the quadrupole effort and delay the start of dipole correctors, sextupoles and trim quads and put measurement of the Optical Klystron on hold. The engineering services contract, which is in the bidding stage, will provide some relief, but it is more expensive and more cumbersome. We are investigating alternative solutions to this potential schedule issue.

* We continued layout of the GX Prototype, refining the details of coil winding configuration to avoid asymmetrical bundle configuration in the beam zone of the magnet face.
* We are evaluating the effect of face angles on the yoke's length with the aim of achieving the correct magnetic length.
* The copper vendor indicated the prototype quantity of custom sized conductor is a viable order.
* We are working on a source for the prototype steel.
* We quickly completed the 3D magnetic model of the existing H style dipole. The model agreed with magnet test data and hence benchmarked the code. We are now constructing the 3D model of the GX dipole.

* Prototype 3-inch quadrupole yoke parts are back from annealing and are being semifinal machined by an outside firm.
* The finishing cut for the yokes above is the first instance of complete CAD?CAM (Computer Aided Design-Computer Aided Machining) at Jefferson Lab. After a number of tries, the digital model of the quadrupole yoke became the basis for the program for yoke machining.

Wiggler (WBS 10):

* We continued layout and details of the Optical Klystron on its strong back with all diagnostics, vacuum pipe, dispersion section and water manifolds
* We stiffened the mount for the dispersion section yokes and, to make alignment easier, inverted the orientation of the adjuster cartridge.

Optics (WBS 11):

* One upgrade and one new copy of GLAD was purchased and initial trials of axicon pair beam compression was modeled.
* Conversion of the 2D Optical Transport drawing into a 3D version was completed. This will allow us to identify and fix potential interferences.


As part of this user run we have been utilizing crews from CEBAF Operations. The operators are being trained in not only the linac setup but also in running the laser so that at some point in the future standard user operations can be taken over by the Machine Control Center. The crews are coming up to speed rapidly and having the extra help is proving to be beneficial.

The FEL has been running very well with the photocathode providing full charge for less than 40% of the available drive light. Efficiency of the FEL at 3 microns has also been exceptional, at least 1.5% in pulse mode or three times specifications. No major glitches have been encountered except for a puzzling asymmetry of the optical mode as measured upstairs. We have replaced and re-measured a series of transport mirrors but the source of this effect has not yet been found. Work over the coming weekend will investigate this effect further. We also plan an attempt at third harmonic lasing.

Competed the installation of X-ray diffractor crystal motion controls and readouts. Continued searching for suitable scintillator for spatial distribution part of xray experiment scheduled for late July. Evaluating the feasibility of IR-X-ray pump-probe part of X-ray experiment.

Updated LSS PLC software in Labs 1 & 4 and re-certified. Upgraded LCW temperature instrumentation to monitor both input and output temperatures of the mirror cooling water.